OpenStudy (anonymous):

Please help A 3.3 kg mass with an initial velocity of v1 = aˆi, where a = 3.47 m/s collides with and sticks to a 0.989 kg mass with an initial velocity of v1 = bˆj, where b = −4.8 m/s. Find the final speed of the composite mass. Answer in units of m/s

OpenStudy (souvik):

it is a perfectly inelastic collision.. linear momentum will be conserved.. let the x component of the final velocity is \(v _{x}\) and the y component of that is \(v _{y}\) in x direction..\[m _{1}*a+m _{2}*0=(m _{1}+m _{2})*v _{x}\] and in y direction..\[m _{1}*0+m _{2}*b=(m _{1}+m _{2})*v _{y}\] now the final velocity will be \[\large{v=\sqrt{v ^{2}_{x}+v ^{2}_{y}}}\]

OpenStudy (anonymous):

I cant see the equations :/

OpenStudy (anonymous):

I've gotten the answer of about 3 and it is wrong

OpenStudy (anonymous):

Also I got 1.56